Seven less than twice a number x is 19. What is the value of 4x?
Answer by Jstrasner(112)

Hey,
So for this one we need to write what we know in numbers and variables:
2x - 7 = 19
To solve this equation, we need to isolate the x. So first we add 7 to both sides to get rid of it on the left side:
2x - 7 + 7 = 19 + 7 => 2x = 26
Now we divide both sides by 2 to get rid of the 2 on the left side:
2x/2 = 26/2 => x = 13
Now we need to solve 4x. Since x = 13, we substitute 13 for x in the equation:
4 (13) = 52
I hope this helps!